# Talk2Radar: Bridging Natural Language with 4D mmWave Radar for 3D Referring Expression Comprehension

### v1

@@ Abstract
Embodied perception is essential for intelligent vehicles and robots, enabling more natural interaction and task execution. However, these advancements currently embrace vision level, rarely focusing on using 3D modeling sensors, which limits the full understanding of surrounding objects with multi-granular characteristics. Recently, as a promising automotive sensor with affordable cost, 4D Millimeter-Wave radar provides denser point clouds than conventional radar and perceives both semantic and physical characteristics of objects, thus enhancing the reliability of perception system. To foster the development of natural language-driven context understanding in radar scenes for 3D grounding, we construct the first dataset, Talk2Radar, which bridges these two modalities for 3D Referring Expression Comprehension. Talk2Radar contains 8,682 referring prompt samples with 20,558 referred objects. Moreover, we propose a novel model, T-RadarNet for 3D REC upon point clouds, achieving state-of-the-art performances on Talk2Radar dataset compared with counterparts, where Deformable-FPN and Gated Graph Fusion are meticulously designed for efficient point cloud feature modeling and cross-modal fusion between radar and text features, respectively. Further, comprehensive experiments are conducted to give a deep insight into radar-based 3D REC. We release our project at https://github.com/GuanRunwei/Talk2Radar.
Index Terms:
3D referring expression comprehension, 3D visual grounding, 4D mmWave radar, multi-modal fusionI Introduction
@@
Millimeter-wave (mmWave) radar, as an all-weather and low-cost perception sensor [1, 2] which can capture objects’ distance, azimuth, velocity, motion directions, reflected power, etc, has been commonly used in perception for autonomous driving [3, 4, 5, 6, 7, 8], robotic navigation [9, 10] and Cooperative Intelligent Transportation Systems (C-ITS) [11]. Recently, to enhance the detection accuracy, 4D radar appears and addresses the limitation of conventional radar’s inability to measure height and significantly increases Point Cloud (PC) density [12, 13]. This allows captured objects to contain more exploitable features for more fruitful scene understanding and better performance in different downstream tasks on land [14, 15, 16, 17, 18, 19] and maritime [20, 21, 22].
@@
Meanwhile, with recent advancements driven by Vision-Language Models (VLMs) [23] in embodied intelligence and human-centric intelligent driving perception systems [24], VLMs and Multi-Modal Large-Language Models (MM-LLMs) enable intelligent vehicles and robots to understand human commands, perceive surroundings, and make corresponding decisions [25]. However, these advancements are primarily limited to the visual domain [26, 27, 28]. While some research focuses on 3D visual grounding, these efforts are primarily aimed at RGB-D and LiDAR [29, 30, 31, 32, 33]. However, for autonomous driving and navigation, we argue that referring prompts cannot only emphasize positional relations, shapes, and categories, but target motion characteristics should be included, which is precisely the strength of radar. Radar is immune to adverse weather and can provide quantified distance, velocity, motion and azimuth while Radar Cross Section (RCS) and PCs provide qualitative semantic features, which enables great flexibility in describing objects, allowing for including both qualitative and quantitative characteristics. With the emergence of high-resolution 4D radar, this unexplored field is expected to be addressed. We aim to pioneer and release the potential of 4D radar for natural language-based multi-modal 3D target localization, precisely, individuals can locate specific objects with the format of 3D bounding box by describing features that 4D radar can perceive.
@@
Based on these, we introduce 3D Referring Expression Comprehension (REC), also called as open-vocabulary 3D object detection [34], that locates specific objects based on textual descriptions, and establish the first radar-based 3D visual grounding dataset, Talk2Radar, and corresponding benchmark. Talk2Radar is built upon the well-known 4D radar dataset, View of Delft (VoD) [35], where each sample contains radar point clouds, LiDAR point clouds, an RGB image, and a textual reference prompt. Moreover, Talk2Radar contains the following two characteristics: (i) A prompt refers to one or more objects, and is more flexible and realistic. (ii) Text features contain target attributes perceivable by radar, while excluding any features perceivable by vision. Moreover, along with Talk2Radar, we propose a novel radar-text fusion model called T-RadarNet for 3D visual grounding. In T-RadarNet, firstly, we propose Deformable-FPN for the enhancement of irregular radar point cloud modeling; secondly, to efficiently fuse the features of radar and text, we devise a module called Gated Graph Fusion (GGF), building upon graph and gating mechanisms for neighborhood feature association and cross-modal feature matching. Notably, GGF can be transplanted to most PC detectors for 3D visual grounding with impressive generalization. We summarize our contributions as follows:
@@
This paper presents the first 4D radar-based 3D REC dataset, Talk2Radar, and its corresponding benchmark for subsequent studies.
@@
A novel 3D REC model called T-RadarNet for radar-based 3D visual grounding, building upon devised Deformable-FPN and GGF, is proposed.
@@
Comprehensive experiments are conducted to analyze and advance 4D radar-based visual grounding, thus also facilitate a holistic understanding in this field.
@@ II Related Works
As a challenging task based on multi-modal learning and 3D geometry, 3D REC aims to provide a textual prompt and locate one or more objects most closely matching it in the form of 3D bounding boxes. (i) Datasets. As Fig. I shows, [36] propose a 2D image-based REC dataset on driving scenarios upon Nuscenes [37], but providing LiDAR data of the same frame. [38] raises a monocular 3D grounding dataset upon KITTI [39]. [40] proposes a 3D grounding benchmark based on multi-camera and multi-frames for autonomous driving. (ii) Methods. [38] proposes a monocular-based 3D REC baseline model called Mono3DVG-TR. [41] raises a model for modeling complex 3D object relation called AFMNet. [31] propose a baseline model for LiDAR-based visual grounding upon Talk2Car dataset. Nevertheless, radar-based 3D REC is still unexplored and lacks full-scale analysis. Further, textual prompts of our Talk2Radar dataset contain more attributes of objects and context with qualitative or numerical description, and provide a realistic referring paradigm.
II-B 3D Object Detection based on Point Cloud
3D object detection is vital for environmental perception for autonomous driving. Point cloud sensors mainly include LiDAR and 4D radar. LiDAR can provide rich geometric and depth features, while 4D radar can capture certain semantic features, motion, velocity, and depth features of objects. Currently, PC-based 3D detectors can be primarily divided into pillar-based and voxel-based. [42, 43, 44] are three fast pillar-based, which project point clouds onto a Bird’s-Eye View (BEV) and then extract features like image processing. [45, 46, 47] are three voxel-based detectors converting point clouds into voxel grids, which are known for simplicity, but they may suffer from low computation efficiency. Additionally, [48, 49] provide an anchor-free paradigm for 3D detection. [50] fuses different representations of point clouds dedicated to radar-based 3D detection. Based on the above, we fully consider two aspects in our T-RadarNet: (i) Extraction of irregular point cloud features, (ii) Dynamic weighting and separation of objects and clutter for high-quality fusion with text features.
@@
| NuPro [40] | Camera | 187,445 | - | color size relation |
| motion location | ||||
| Talk2Car [36] | Camera | 10,519 | 11.01 | color size |
| location relation | ||||
| M3DRef [38] | Camera | 8,228 | 53.24 | size relation |
| location color | ||||
| Talk2Radar | 4D radar, | 20,558 | 14.30 | size motion location |
| LiDAR | relation velocity depth |
@@ III-A Data Collection and Annotation Process
Thanks to the renowned VoD dataset [35] in the field of radar perception, which is specifically designed for autonomous driving perception, equipped with the ZF FRGen21 4D mmWave radar, as well as a LiDAR and a stereo camera. We can engineer textual prompts upon finely annotated 3D bounding boxes of various road objects. Fig. 1 shows the samples while Fig. 2 presents the whole annotation process.
(1) Select Candidate Object(s). First, based on the developed Graphical User Interface (GUI) for annotation, we present the image with projected 3D bounding boxes for easy observation. Then, a specialist (annotator) with experience in the traffic clicks and selects the candidate boxes in the image that want to describe. The corresponding 3D bounding box annotation will be saved. Here, the annotator prioritizes selecting objects within the radar’s Field of View (FoV).
(2) Check Annotations and Write Prompts. Based on chosen 3D bounding box(es), the annotator first check whether the bounding box(es) have errors on position or category. During annotation, we also attach compensated vector velocity and plane depth next to the corresponding PC for reference. Considering the detection error and object size, we generally describe the object depth as a range rather than a fixed value. For individual objects, annotators consider attributes such as category, spatial position relative to the ego vehicle, distance, velocity, motion trend, size, etc., but do not include color, and other information that radar cannot perceive. If multi-objects are selected, their spatial relationships are also considered.
(3) Enrich Annotation and Correct Text Bias. To avoid subjectivity and errors in text prompts, we implement a collaborative review and revision process involving two additional annotators after the initial annotation. This process serves to increase the diversity of descriptions to test the robustness and generalization of the model. Additionally, it allows for the discussion of subjective errors made by the initial annotator, followed by corrections upon reaching a consensus.
@@ III-B Dataset Statistics
We elucidate the quantitative statistics from the aspects of textual prompts and referent objects.
Textual Prompts. As Fig. 3 (a) presents, firstly, high-frequency words in the word cloud include both qualitative and quantitative descriptions, where rich mixed vocabulary benefits from the 4D radar’s multi-dimensional perception of objects, capturing quantitative features such as target velocity, motion direction, depth, etc., while also perceiving some degree of semantic features. This enables more detailed descriptions of specific objects and allows for fine-grained target clustering and filtering based on individual or partial target attributes. Secondly, as Fig. 3 (c) shows, the prompt length distribution in Talk2Radar is broad, whose average value is 14.30 (Table I, AvgExpr), allowing the model to learn a wide range of sentence patterns. This also increases the task difficulty, as the model must handle text prompts of varying patterns, structures, and lengths, correctly understand their hidden semantic representations, and query complex point cloud scenes.
Referent Objects. Firstly, as Table I presents, Talk2Radar contains 20,558 referent objects totally. Secondly, Table II shows the distribution of target quantities across 12 categories and the average PC quantity for each target category, including radar (1, 3 and 5 frames accumulated) and LiDAR. Car, Pedestrian and Cyclist are three primary referent categories in Talk2Radar. Evidently, although 4D radar has improved resolution, the gap in PC density compared to LiDAR still exists. Thirdly, as Fig. 3 (b) shows, in Talk2Radar, the referent object number of each sample range from 1 to 11, presenting a challenge for understanding complex scenes. Specifically, the model needs to adaptively filter out noise words in prompts and focus on the key characteristics of the referent object(s), while performing multiple Region-Of-Interest (ROI) queries in irregular point cloud contexts.
@@ III-C Metrics and Subset Settings
Metrics. We adopt Average Precision (AP) and Average Orientation Similarity (AOS), including 3D bounding box mAP and mAOS on the entire annotated and driving corridor area, consistent with VoD for comprehensive evaluation.
Subset Settings. We divide Talk2Radar into three subsets, sharing same IDs with VoD. The number of subsets for training, validation and test are 5139, 1296 and 2247, respectively. Since the test set annotation of VoD is not publicly available, models are evaluated on the validation set in this paper.
@@ IV-A Overall Pipeline
Fig. 4 shows the detailed architecture of T-RadarNet. Given a frame of radar PC and a textual prompt as input, T-RadarNet will give the 3D grounding prediction of the specific object(s) upon the guidance of natural language. Firstly, for the primary PC feature extraction and modeling, considering efficient computation and scalability, we adopt the pillar encoder for 3D PC basic representation as a 2D pseudo image. Then, the concise SECOND backbone [46] is to extract three-stage radar PC features with multi-scales. Secondly, we adopt transformers-based encoders (e.g. ALBERT) [51] for dynamic context representation of textual prompt. Thirdly, given the interference in radar PC context from multi-path clutter, aligning and fusing dense textual semantic features with sparse point clouds pose significant challenges. To address this, we design a graph-based strategy called Gated Graph Fusion (GGF), which associates ROI in the radar feature space and applies point-wise cross-modal gating between radar and text, yielding the most matched point cloud regions aligned with the text. Then, three-stage PC features and textual feature are fed into GGF-FPN, which consists of GGF modules for cross-modal fusion and outputs three-scale text-conditional radar features . Fourthly, given the irregularity and sparsity of radar PC, conventional convolutions struggle to finely model them. Hence, we propose Deformable-FPN based on deformable convolutions [52], which provide high-quality PC features queried by the textual prompt to the grounding head for localization. Lastly, leveraging the efficiency of anchor-free detectors, we adopt the center-based detection head [48] and employ the deformable convolution-based separable localization head for text-referenced point cloud object localization.
@@ IV-B Gated Graph Fusion
For one-stage deep radar PC feature , we first construct a graph-based map to aggregate neighbour features and associate the potential object areas. We implement the graph convolution as Eq. (1) presents:
where and are two learnable weights for feature aggregation and update. Exactly, the aggregation operation calculates the representation of the current node by aggregating features of neighbouring nodes, which is presented in Eq. (2):
where is the set of neighbour nodes of . Considering the efficiency of graph feature aggregation, we employ Max-Relative Graph Convolution [53], which is shown in Eq. (3) and (4) as below:
The above constructs the radar graph feature in one stage of three.
For the textual feature , we first adopt max-pooling along the token dimension to extract abstract high-level semantic information. Then, to align with the radar graph feature , a linear transformation is exerted on it and we obtain the abstract textual feature . Further, to effectively suppress radar noise points and focus on object ROI, we adopt a cross-modal gating strategy. Specifically, we obtain the weights of the linearly transformed textual features through Sigmoid activation and multiply them point-wise with the radar graph feature . Finally, the feature is enhanced by addition with by a residual path. The process is shown in Eq. (5) and (6):
where and denote point-wise multiplication and Sigmoid.
@@
| PointPillars (SFPN) | Radar5 | ALBERT | HDP | 18.92 | 9.79 | 12.47 | 13.73 | 12.91 | 39.20 | 10.25 | 14.93 | 21.46 | 20.19 |
| SECOND (SFPN) | Radar5 | ALBERT | HDP | 17.70 | 7.67 | 10.58 | 11.98 | 11.06 | 38.17 | 8.62 | 14.95 | 20.58 | 19.77 |
@@
| CenterPoint (Pillar-SFPN) | Radar5 | ALBERT | HDP | 18.98 | 5.30 | 14.96 | 13.08 | 12.20 | 40.53 | 8.57 | 15.66 | 21.59 | 20.25 |
@@
| CenterPoint (Pillar-SFPN) | Radar5 | ALBERT | MHCA | 5.21 | 4.57 | 5.13 | 4.97 | 3.12 | 12.70 | 4.07 | 7.70 | 8.16 | 7.51 |
| MSSG [31] | Radar5 | GRU | - | 12.53 | 5.08 | 8.47 | 8.69 | 7.03 | 18.93 | 7.88 | 9.40 | 12.07 | 11.67 |
@@
| T-RadarNet | Radar5 | ALBERT | GGF | 24.68 | 9.71 | 15.74 | 16.71 | 14.88 | 42.58 | 10.13 | 17.82 | 23.51 | 22.37 |
@@
| T-RadarNet | LiDAR | ALBERT | GGF | 24.91 | 12.74 | 18.67 | 18.77 | 17.20 | 48.98 | 14.69 | 27.24 | 30.30 | 29.89 |

### v3

@@ Abstract
Embodied perception is essential for intelligent vehicles and robots in interactive environmental understanding. However, these advancements primarily focus on vision, with limited attention given to using 3D modeling sensors, restricting a comprehensive understanding of objects in response to prompts containing qualitative and quantitative queries. Recently, as a promising automotive sensor with affordable cost, 4D millimeter-wave radars provide denser point clouds than conventional radars and perceive both semantic and physical characteristics of objects, thereby enhancing the reliability of perception systems. To foster the development of natural language-driven context understanding in radar scenes for 3D visual grounding, we construct the first dataset, Talk2Radar, which bridges these two modalities for 3D Referring Expression Comprehension (REC). Talk2Radar contains 8,682 referring prompt samples with 20,558 referred objects. Moreover, we propose a novel model, T-RadarNet, for 3D REC on point clouds, achieving State-Of-The-Art (SOTA) performance on the Talk2Radar dataset compared to counterparts. Deformable-FPN and Gated Graph Fusion are meticulously designed for efficient point cloud feature modeling and cross-modal fusion between radar and text features, respectively. Comprehensive experiments provide deep insights into radar-based 3D REC. We release our project at https://github.com/GuanRunwei/Talk2Radar.
I Introduction
@@
Millimeter-wave (mmWave) radar, as an all-weather and low-cost perception sensor [1, 2] which can capture objects’ distance, azimuth, velocity, motion directions, reflected power, etc., has been widely used in perception for autonomous driving [3, 4, 5, 6, 7, 8], robotic navigation [9, 10] and Cooperative Intelligent Transportation Systems (C-ITS) [11]. Recently, to enhance detection accuracy, 4D radar emerges and addresses the limitation of conventional radars’ inability to measure height and significantly increases Point Cloud (PC) density [12, 13], allowing objects to contain more exploitable features for fruitful scene understanding and better performance in downstream tasks on land [14, 15, 16, 17, 18, 19, 20, 21] and maritime [22, 23, 24, 25, 26].
@@
Meanwhile, recent advancements driven by Vision-Language Models (VLMs) [27] in embodied intelligence and human-centric intelligent driving perception systems [28] enable intelligent vehicles and robots to understand human commands, perceive surroundings, and make decisions [29]. However, these advancements are primarily confined to the visual domain [30, 31, 32]. While some research has focused on 3D visual grounding, these efforts are predominantly aimed at RGB-D and LiDAR [33, 34, 35, 36, 37]. For autonomous driving and navigation, we argue that referring prompts should not only emphasize positional relations, shapes, and categories but also include object motion characteristics, which is precisely the strength of radar. Radar is immune to adverse weather conditions and can provide quantified measurements of distance, velocity, motion, and azimuth, while Radar Cross Section (RCS) and Point Clouds (PCs) offer qualitative semantic features. This dual capability allows for a comprehensive description of objects, incorporating both qualitative and quantitative characteristics. With the emergence of high-resolution 4D radar, this unexplored field is poised to be addressed. Unlike VLM-based environmental understanding, which can leverage web-scale training data, radar data depend on real-world scenarios and physical models [38]. Currently, there are no relevant datasets or research focusing on this area. Hence, we aim to pioneer and unlock the potential of 4D radar for natural language-based multi-modal 3D object localization. Specifically, this enables individuals to locate specific objects using 3D bounding boxes by describing features perceivable by 4D radar.
@@
Based on these foundations, we introduce 3D Referring Expression Comprehension (REC), also known as open-vocabulary 3D object detection or 3D visual grounding [39], which locates specific objects based on textual descriptions. To this end, we establish the first radar-based 3D visual grounding dataset, Talk2Radar, and its corresponding benchmark. Talk2Radar is built upon the well-known 4D radar dataset, View of Delft (VoD) [40]. Each sample in Talk2Radar contains radar PCs, LiDAR PCs, an RGB image, and a textual reference prompt. The dataset possesses the following two key characteristics: (i) A prompt can refer to one or more objects, making it more flexible and realistic; (ii) The text features contain object attributes perceivable by radar, while excluding any features perceivable by vision. Besides the dataset, we propose a novel radar-text fusion model called T-RadarNet for 3D visual grounding. Within T-RadarNet, firstly, we introduce Deformable-FPN to enhance the modeling of irregular radar PCs. Secondly, to efficiently fuse the features of radar and text, we devise a module named Gated Graph Fusion (GGF). GGF leverages graph and gating mechanisms for neighborhood feature association and cross-modal feature matching. Notably, GGF can be integrated into most PC detectors for 3D visual grounding, demonstrating impressive generalization capabilities.
The contributions of this paper are summarized as follows:
@@
We present Talk2Radar, the first 4D radar-based 3D REC dataset for terrestrial autonomous driving, along with its corresponding benchmark for subsequent studies.
@@
For 4D radar-based 3D visual grounding, we propose a novel 3D REC model named T-RadarNet, by incorporating the devised Deformable-FPN and GGF.
@@
Comprehensive experiments have been conducted to analyze and enhance 4D radar-based 3D visual grounding, to promote a thorough understanding in this field.
@@ II Related Works
As a challenging task based on multi-modal learning and 3D geometry, 3D REC aims to provide a textual prompt and locate one or more objects most closely matching it in the form of 3D bounding boxes. (i) Datasets. As Table. I shows, Talk2Car [41] is a 2D image-based REC dataset on driving scenarios upon Nuscenes [42], but providing LiDAR data of the same frame. M3DRef [43] is a monocular 3D grounding dataset upon KITTI [44]. NuPro [45] is a 3D grounding benchmark based on multi-camera and multi-frames for autonomous driving. (ii) Methods. [43] proposes a monocular-based 3D REC baseline model called Mono3DVG-TR. [46] raises a model called AFMNet for modeling complex 3D object relation. [35] propose a baseline model for LiDAR-based visual grounding upon Talk2Car dataset. Nevertheless, radar-based 3D REC is still unexplored and lacks full-scale analysis. Moreover, textual prompts of our Talk2Radar dataset contain more attributes of objects and context with qualitative or numerical description, and provide a realistic referring paradigm.
II-B 3D Object Detection with Point Cloud
3D object detection is vital for environmental perception for autonomous driving. PC sensors mainly include LiDAR and 4D radar. LiDAR can provide rich geometric and depth features [47, 48], while 4D radar can capture certain semantic features, motion, velocity, and depth features of objects [49, 50, 51]. Currently, PC-based 3D detectors can be primarily divided into pillar-based and voxel-based. [52, 53, 54] are three fast pillar-based, which project PCs onto a Bird’s-Eye View (BEV) and then extract features like image processing. [55, 56, 57] are three voxel-based detectors converting PCs into voxel grids, which are known for simplicity, but they may suffer from low computation efficiency. Additionally, [58, 59] provide an anchor-free paradigm for 3D detection. [60] fuses different representations of PCs dedicated to radar-based 3D detection. Based on the above, we fully consider two aspects in our T-RadarNet: (i) Extraction of irregular PC features, (ii) Dynamic weighting and separation of objects and clutter for high-quality fusion with text features.
@@
| NuPro [45] | Camera | 187,445 | - | color, size, relation |
| motion, location | ||||
| Talk2Car [41] | Camera, | 10,519 | 11.01 | color, size, |
| LiDAR | location, relation | |||
| M3DRef [43] | Camera | 8,228 | 53.24 | size, relation, |
| location, color | ||||
| Talk2Radar | 4D radar, | 20,558 | 14.30 | size, motion, location, |
| LiDAR | relation, velocity, depth |
@@ III-A Data Collection and Annotation Process
Thanks to the renowned VoD dataset [40] in the field of radar perception, which is specifically designed for autonomous driving perception and equipped with the ZF FRGen21 4D mmWave radar, as well as a LiDAR and a stereo camera, we can engineer textual prompts based on finely annotated 3D bounding boxes of various road objects. Fig. 1 shows some samples, while Fig. 2 presents the entire annotation process.
(1) Select Candidate Object(s): First, utilizing the developed Graphical User Interface for annotation, we present the image with projected 3D bounding boxes for easy observation. Then, experienced annotators select the candidate boxes in the image that need to be described. The corresponding 3D bounding box annotation is saved. The annotator prioritizes selecting objects within the radar’s Field of View (FoV).
(2) Check Annotations and Write Prompts: Based on the chosen 3D bounding box(es), the annotator first checks for any errors in the position or category of bounding boxes. During annotation, we also attach compensated vector velocity and plane depth next to the corresponding PC for reference. Considering detection errors and object size, we generally describe the object depth as a range rather than a fixed value. For individual objects, annotators consider attributes such as category, spatial position relative to the ego vehicle, distance, velocity, motion trend, and size, but exclude color and other information that radar cannot perceive. If multiple objects are selected, their spatial relationships are also considered.
(3) Enrich Annotation and Correct Text Bias: To avoid subjectivity and errors in text prompts, we implement a collaborative review and revision process involving two additional annotators after the initial annotation. This process serves to increase the diversity of descriptions to test the model’s robustness and generalization. Additionally, it allows for the discussion of subjective errors made by the initial annotator, followed by corrections upon reaching a consensus.
@@ III-B Dataset Statistics
We present quantitative statistics from the perspectives of textual prompts and referent objects.
Textual Prompts: As illustrated in Fig. 3(a), firstly, high-frequency words in the word cloud include both qualitative and quantitative descriptions. This rich vocabulary benefits from the 4D radar’s multi-dimensional perception capabilities, capturing quantitative features such as object velocity, motion direction, and depth, while also perceiving some degree of semantic features. This enables more detailed descriptions of specific objects and allows for fine-grained object clustering and filtering based on individual or partial object attributes. Secondly, as shown in Fig. 3(c), the prompt length distribution in Talk2Radar is broad, with an average value of 14.30 (Table I, AvgExpr). This variety allows the model to learn a wide range of sentence patterns, increasing task complexity as the model must handle text prompts of varying patterns, structures, and lengths, correctly understanding their semantic representations to query complex PC scenes.
Referent Objects: Firstly, as presented in Table I, Talk2Radar contains a total of 20,558 referent objects. Secondly, Table II shows the distribution of object quantities across 12 categories and the average PC quantity for each category, including radar (1, 3, and 5 frames accumulated) and LiDAR. Cars, pedestrians, and cyclists are the three primary referent categories in Talk2Radar. Despite the improved resolution of 4D radar, there remains a gap in PC density compared to LiDAR. Thirdly, as shown in Fig. 3(b), the number of referent objects per sample in Talk2Radar ranges from 1 to 11, presenting a challenge for understanding complex scenes. Specifically, the model must adaptively filter out noise words in prompts, focus on the key characteristics of the referent object(s), and perform multiple Region-Of-Interest (ROI) queries in irregular PC contexts.
@@ III-C Metrics and Subset Settings
Metrics: We adopt Average Precision (AP) and Average Orientation Similarity (AOS), including 3D bounding box mAP and mAOS on the entire annotated and driving corridor area, consistent with VoD for comprehensive evaluation.
Subset Settings: We divide Talk2Radar into three subsets, sharing same IDs with VoD. The number of subsets for training, validation and test are 5139, 1296 and 2247, respectively. Since the test set annotation of VoD is not publicly available, models are evaluated on the validation set in this paper.
@@ IV-A Overall Pipeline
Fig. 4 illustrates the detailed architecture of T-RadarNet. Given a frame of radar PCs and a textual prompt as input, T-RadarNet provides the 3D grounding prediction of the specified object(s) guided by natural language. Firstly, for primary PC feature extraction, considering computational efficiency and scalability, we adopt a pillar encoder for the basic representation of 3D PC as a 2D pseudo image. Subsequently, the concise SECOND backbone [56] is used to extract three-stage radar PC features at multiple scales. Secondly, we use transformer-based encoders (e.g., ALBERT) [61] for dynamic context representation of the textual prompt. Thirdly, given the interference in radar PC context from multi-path clutter, aligning and fusing dense textual semantic features with sparse PCs pose significant challenges. To address this, we design a graph-based strategy called Gated Graph Fusion (GGF), which associates ROI in the radar feature space and applies point-wise cross-modal gating between radar and text, yielding the most matched PC regions aligned with the text. The three-stage PC features and textual feature are then fed into GGF-FPN, which consists of GGF modules for cross-modal fusion, outputting three-scale text-conditional radar features . Fourthly, given the irregularity and sparsity of radar PC, conventional convolutions struggle to interact with them. Hence, we propose Deformable-FPN based on deformable convolutions [62], which provides high-quality PC features queried by the textual prompt to the grounding head for localization. Lastly, considering the efficiency of anchor-free detectors, we adopt the center-based detection head [58] and employ the deformable convolution-based separable localization head for text-referenced 3D detection.
@@ IV-B Gated Graph Fusion
For one-stage deep radar feature , we first construct a graph map to aggregate neighbour features and associate potential object areas, where denotes a set of unordered nodes initialized upon [63]. We implement the graph convolution as:
where and are two learnable weights for feature aggregation and update. Specifically, the aggregation and update operations calculate the representation of the current node by aggregating features of neighbouring nodes, i.e.,
where is the set of neighbour nodes of . Considering the efficiency of graph feature aggregation, we first adopt Max-Relative Graph Convolution (MRConv4d) [63]. To enrich context, we utilize a residual path with concatenation to construct radar graph with neighourhood aggregated features:
Based on the above, we construct the graph-based deep radar feature .
For the textual feature , we first adopt max-pooling along the token dimension to extract abstract high-level semantic information. Then, to align with the radar graph feature , a linear transformation is exerted on it and we obtain the abstract textual feature . Further, to effectively suppress noise and focus on object ROI, we adopt a cross-modal gating strategy. Exactly, we obtain the weights of the linearly transformed textual features through Sigmoid activation and multiply them point-wise with the radar graph feature . Finally, the feature is enhanced by addition with by a residual path. The process is shown below:
where and denote the point-wise multiplication and Sigmoid function, respectively.
@@
| PointPillars (SFPN) | Radar5 | ALBERT [64] | HDP | 18.92 | 9.79 | 12.47 | 13.73 | 12.91 | 39.20 | 10.25 | 14.93 | 21.46 | 20.19 |
@@
| CenterPoint (Voxel-SFPN) | Radar5 | ALBERT | HDP | 18.98 | 5.30 | 14.96 | 13.08 | 12.20 | 40.53 | 8.57 | 15.66 | 21.59 | 20.25 |
@@
| CenterPoint (Voxel-SFPN) | Radar5 | ALBERT | MHCA | 5.21 | 4.57 | 5.13 | 4.97 | 3.12 | 12.70 | 4.07 | 7.70 | 8.16 | 7.51 |
| MSSG [35] | Radar5 | GRU [65] | - | 12.53 | 5.08 | 8.47 | 8.69 | 7.03 | 18.93 | 7.88 | 9.40 | 12.07 | 11.67 |
@@
| MSSG [35] | Radar5 | ALBERT | - | 16.03 | 5.86 | 10.57 | 10.82 | 8.96 | 25.79 | 8.69 | 12.55 | 15.68 | 14.12 |
| AFMNet [46] | Radar5 | ALBERT | - | 16.31 | 6.80 | 10.35 | 11.15 | 9.46 | 26.82 | 8.71 | 12.45 | 15.99 | 14.18 |
| EDA [66] | Radar5 | RoBERTa [67] | - | 13.23 | 6.60 | 8.63 | 9.49 | 8.93 | 23.55 | 8.80 | 11.95 | 14.77 | 13.07 |
| T-RadarNet (Ours) | Radar5 | ALBERT | GGF | 24.68 | 9.71 | 15.74 | 16.71 | 14.88 | 42.58 | 10.13 | 17.82 | 23.51 | 22.37 |
@@
| MSSG [35] | LiDAR | ALBERT | - | 18.19 | 7.66 | 11.63 | 12.49 | 10.91 | 29.61 | 10.98 | 14.66 | 18.42 | 16.23 |
| AFMNet [46] | LiDAR | ALBERT | - | 19.50 | 7.92 | 13.56 | 13.66 | 12.18 | 31.68 | 9.23 | 18.90 | 19.94 | 17.59 |
| EDA [66] | LiDAR | RoBERTa | - | 16.10 | 6.91 | 12.88 | 11.96 | 10.10 | 25.10 | 9.28 | 15.73 | 16.70 | 14.91 |
| T-RadarNet (Ours) | LiDAR | ALBERT | GGF | 24.91 | 12.74 | 18.67 | 18.77 | 17.20 | 48.98 | 14.69 | 27.24 | 30.30 | 29.89 |